# Multi-Layered Reasoning from a Single Viewpoint for Learning See-Through Grasping

### v4

@@ 4.2.2 Multi-Modal Visual Perception of the Occluded Scene via In-Finger Vision
An intuitive understanding of the second research problem is whether we can utilize in-finger vision for environmental perception, which is inherently multi-modal, thereby eliminating the need for an external camera in modern robot workstations.
@@
Using in-finger vision for environmental perception is counterintuitive. Having an Eye camera fixed to the robot base is generally perceived as bio-inspired by humans or most living creatures, where visual organs are on the body for environmental perception. This is also a feature shared by most robotic systems. After masked segmentation of the See-Thru-Network, the original image becomes fragmented. Still, it could be alternatively repaired using inpainting, a highly active field of research in computer vision and widely adopted practice throughout history. A direct result of inpainting the in-finger vision is the reconstruction of the scene, producing visual perceptual capabilities almost precisely as a typical Eye camera. It should be noted that the inpainting technique is computationally intensive, requiring high-performing hardware and long processing time, neither of which is ideal for robotic applications. In this study, we added an extra head-up motion during motion planning to collect an extra 1-second video clip for inpainting as a trade-off.
@@
In the sections above, we demonstrated object detection using inpainted images from in-finger vision. Here, we showcase the performance of achieving other modalities, such as depth sensing and scene segmentation, using the inpainted images, which are also widely used in modern robotics. In principle, robotists can reuse the inpainted images for any desirable extra modalities as needed, just as in a computer vision problem, making the VBSeeThruP architecture a generic solution for robotics, where constraints in hardware cost and computational power are significant considerations.
@@
We estimated monocular depth based on the inpainted scene as proof of concept. We compared the reconstructed scene depths from both the in-finger camera and a camera with an unobstructed view of the robot’s working space. In this experiment, ten objects with various shapes and colors (milk powder can, tape, rubber ball, strawberry, glue, plum, glue stick, power drill with a packaging box, and drill toolkit) were placed on the tabletop. The robotic arm swept across the tabletop from left to right and returned to the starting point with a fixed orientation while the in-finger camera inspected the objects. We recorded both the in-finger and unobstructed videos by removing the SPN from the gripper. The former was further inpainted by the E2FGVI algorithm. The inpainted and unobstructed videos are then fed to the Depth Anything V2 model [71] to obtain the depth estimations. The pixel-wise estimations provide relative depth values, offering helpful information on the relative poses of objects on the table. Hence, instead of comparing the actual values of the depth map, we adopt the structural similarity index (SSI) to evaluate the similarity between the depth map reconstructed from the inpainted view and that from the unobstructed view.
As shown in Fig. 11A, the index varies between 0.71 and 0.95 with an average of 0.87. Figure 11B-D shows three snapshots of in-finger vision, the depth map estimated from the inpainted view, and the SSI map. Low SSI values prominently occur at the objects’ edges, where depth estimation is inherently more challenging.
@@
Furthermore, we explored the feasibility of multi-object segmentation on the inpainted scene using the latest Segment Anything Model 2 (SAM2) [72]. We tested on the same scene described in the depth estimation task.
As shown in Fig. 12A, segmentation on the inpainted view achieves satisfactory performance with J measurement ranging between 0.64 and 0.87, which is 16% lower than the performance of segmentation on the unobstructed view on average. An interesting observation occurred in the segmentation of object 9 (a power drill with a packaging box), where the inpainted view performed better than the unobstructed view. Further frame-by-frame examination shows that the worst segmentation in the unobstructed view was caused by the intense surface reflection of the drill’s packaging box, and in some frames, SAM2 failed to segment object 9. On the contrary, SAM2 successfully tracked object nine in the inpainted view. The snapshots of the unobstructed and in-finger vision and segmentations of the inpainted scene at two different times are shown in Figs. 12C&D.
Multi-Modal & Markerless VBSeeThruP
An intuitive understanding of the third research problem is whether we can utilize in-finger vision with a soft finger to achieve high-performing, multi-modal perception in a markerless manner. Similar capabilities have been widely researched in the past literature, but the involvement of artificial markers remains a challenge among the current solutions [73, 74, 75, 76]. In this study, a by-product while addressing the first research problem is the real-time, pixel-wise segmentation of the See-Thru-Network’s mask throughout the implementation of the VBSeeThruP architecture. Visual features from this mask provide a robust sensory input to infer FT information once a calibration dataset is obtained against commercial FT sensors for model training.
While it is reasonable to use simpler neural networks, we adopted a generative approach in this study and utilized the Supervised Variational Autoencoder (SVAE) to address this problem. The latent variables obtained from SVAE are a powerful tool for inferring various sensory modalities, requiring minimal disruption to the overall computation. Besides the 6D FT, we demonstrate a simple implementation to infer the contact event on the soft finger. Recent research also shows promising results in proprioceptive shape reconstruction [17], contact point estimation [34], and tactile perception using in-finger vision [18] with the soft polyhedral network used in this study, which could be further implemented using the deformation mask tracked in this study.
@@
In previous work where the complex deformations of SPN are physically encoded into a much lower 6D space of AruCo Marker pose, the contact location information has been lost [18]. With a much higher dimension of 32 in the SVAE framework, we hypothesize that such information could be maintained in the latent space to a certain degree. A preliminary test was conducted by categorizing the training samples into three contact events: no contact, contact at the finger’s tip, and the middle region. We trained a contact event classifier using the SVAE framework shown in Fig. 7, which can be formulated by
Figure 13 shows the confusion matrix of the classifier evaluated on the validation dataset of 8,000 samples. The accuracies for all three contact events are above 97%. Considering the classifier only takes the SPN’s mask as input, which filters out all the see-through visual information, the results confirm our hypothesis that the soft touch represented by the latent vectors contains information such as force/torque and contact location. A real-time demonstration is included in the supplementary video. The binary classes of the contact location are partly limited by the data collection process, where the precise locations were not recorded. The granularity could be further improved by using denser labels or by converting it into a regression task.
@@ 5 Conclusion, Limitations, and Future Directions
Inspired by sensory substitution from biological agents, this study presents a Vision-based See-Through Perception (VBSeeThruP) approach to reason among different image layers for visual- and contact-based information from a single camera input. By placing a See-Thru-Network in front of a single in-finger camera, two layers of information can be obtained: one at a closer range for the soft interface, awaiting contact interactions, and another about the scene at a distance but occluded. Contact-based perception is substituted markerlessly by reasoning about the masked deformation, facilitating reactive see-through grasping.
@@
Several directions remain open. Promising avenues include underwater perception and manipulation [77], embodied robot learning with multi-modal feedback that integrates SPN’s omni-directional adaptability and precise 6D force/torque sensing [78], and extending the single-camera approach to stereo vision for improved 3D reconstruction or 6D pose estimation.
@@ Acknowledgement
This work was partly supported by the National Natural Science Foundation of China under Grants 62473189 and 62206119, as well as the Shenzhen Long-Term Support for Higher Education at SUSTech under Grant 20231115141649002.
Author Fang Wan would like to thank Prof. Roberto Calandra and all organizers and participants of the 2024 Touch Sensing and Processing Summer School in Dresden for their valuable discussions and constructive feedback on this work.
Supplementary Materials
Please refer to this link for supplementary videos of this paper.

### v5

@@ 4.2.2 Multimodal Visual Perception of the Occluded Scene via In-Finger Vision
An intuitive understanding of the second research problem is whether we can utilize in-finger vision for environmental perception, which is inherently multimodal, thereby eliminating the need for an external camera in modern robot workstations.
@@
Using in-finger vision for environmental perception is counterintuitive. Having an Eye camera mounted on the robot base is generally perceived as bio-inspired by humans or other living creatures, where visual organs are located on the body for environmental perception. This is also a feature shared by most robotic systems. After masked segmentation of the See-Thru-Network, the original image becomes fragmented. Still, it could be alternatively repaired using inpainting, a highly active field of research in computer vision and widely adopted practice throughout history. A direct result of inpainting the in-finger vision is the reconstruction of the scene, producing visual perceptual capabilities almost as good as those of a typical Eye camera. It should be noted that the inpainting technique is computationally intensive, requiring high-performing hardware and long processing time, neither of which is ideal for robotic applications. In this study, we added an extra head-up motion during motion planning to collect an extra 1-second video clip for inpainting as a trade-off.
@@
In the sections above, we demonstrated object detection using in-finger vision inpainted images. Here, we showcase the performance of achieving other modalities, such as depth sensing and scene segmentation, using the inpainted images, which are also widely used in modern robotics. In principle, robotists can reuse the inpainted images for any desired extra modalities, just as in a computer vision problem, making the VBSeeThruP architecture a generic solution for robotics, where constraints on hardware cost and computational power are significant considerations.
@@
We estimated monocular depth from the inpainted scene as a proof-of-concept. We compared the reconstructed scene depths from both the in-finger camera and a camera with an unobstructed view of the robot’s working space. In this experiment, ten objects with various shapes and colors (milk powder can, tape, rubber ball, strawberry, glue, plum, glue stick, power drill with a packaging box, and drill toolkit) were placed on the tabletop. The robotic arm swept across the tabletop from left to right and returned to the starting point with a fixed orientation while the in-finger camera inspected the objects. We recorded both the in-finger and unobstructed videos by removing the SPN from the gripper. The former was further inpainted by the E2FGVI algorithm. The inpainted and unobstructed videos are then fed to the Depth Anything V2 model (Yang et al., 2024) to obtain the depth estimations. Pixel-wise estimations provide relative depth values, offering helpful information about the relative poses of objects on the table. Hence, instead of comparing the actual depth map values, we use the structural similarity index (SSI) to evaluate the similarity between the depth map reconstructed from the inpainted view and that from the unobstructed view.
As shown in Fig. 11A, the index varies between 0.71 and 0.95 with an average of 0.87. Figure 11B-D shows three snapshots of in-finger vision, the depth map estimated from the inpainted view, and the SSI map. Low SSI values are most prominent at the objects’ edges, where depth estimation is inherently more challenging.
@@
Furthermore, we explored the feasibility of multi-object segmentation on the inpainted scene using the latest Segment Anything Model 2 (SAM2) (Ravi et al., 2024). We tested on the same scene described in the depth estimation task.
As shown in Fig. 12A, segmentation on the inpainted view achieves satisfactory performance with J measurement ranging between 0.64 and 0.87, which is 16% lower than the performance of segmentation on the unobstructed view on average. An interesting observation emerged during the segmentation of object 9 (a power drill with a packaging box): the inpainted view performed better than the unobstructed view. Further frame-by-frame examination shows that the worst segmentation in the unobstructed view was caused by the intense surface reflection of the drill’s packaging box, and in some frames, SAM2 failed to segment object 9. On the contrary, SAM2 successfully tracked object nine in the inpainted view. The snapshots of the unobstructed and in-finger vision and segmentations of the inpainted scene at two different times are shown in Figs. 12C&D.
Multimodal & Markerless VBSeeThruP
An intuitive understanding of the third research problem is whether we can utilize in-finger vision with a soft finger to achieve high-performing, multimodal perception in a markerless manner. Similar capabilities have been widely studied in the past literature, but the use of artificial markers remains a challenge among current solutions (Yamaguchi and Atkeson, 2016; Lambeta et al., 2020; Lepora, 2021; Fan et al., 2024). In this study, a by-product while addressing the first research problem is the real-time, pixel-wise segmentation of the See-Thru-Network’s mask throughout the implementation of the VBSeeThruP architecture. Visual features from this mask provide robust sensory input for inferring FT information, once a calibration dataset is obtained against commercial FT sensors for model training.
While simpler neural networks are reasonable, we adopted a generative approach in this study and used the Supervised Variational Autoencoder (SVAE) to address this problem. The latent variables obtained from SVAE are a powerful tool for inferring various sensory modalities, requiring minimal disruption to the overall computation. Besides the 6D FT, we demonstrate a simple method for inferring the contact event on the soft finger. Recent research also shows promising results in proprioceptive shape reconstruction (Guo et al., 2024a), contact point estimation (Guo et al., 2024b), and tactile perception using in-finger vision (Liu et al., 2024b) with the soft polyhedral network used in this study, which could be further implemented using the deformation mask tracked in this study.
@@
In previous work where the complex deformations of SPN are physically encoded into a much lower 6D space of AruCo Marker pose, the contact location information has been lost (Liu et al., 2024b). With a much higher dimension of 32 in the SVAE framework, we hypothesize that such information could be maintained in the latent space to a certain degree. A preliminary test was conducted by categorizing the training samples into three contact events: no contact, contact at the finger tip, and contact in the middle region. We trained a contact event classifier using the SVAE framework shown in Fig. 7, which can be formulated by
Figure 13 shows the confusion matrix of the classifier evaluated on the validation dataset of 8,000 samples. The accuracies for all three contact events are above 97%. Given that the classifier only takes the SPN’s mask as input, which filters out all see-through visual information, the results confirm our hypothesis that the soft touch represented by the latent vectors contains information such as force/torque and contact location. A real-time demonstration is included in the supplementary video. The binary classification of contact locations is partly constrained by the data-collection process, which did not record precise locations. The granularity could be further improved by using denser labels or by converting it into a regression task.
@@ 5 Conclusion, Limitations, and Future Directions
Inspired by sensory substitution from biological agents, this study presents a Vision-based See-Through Perception (VBSeeThruP) approach that reasons across different image layers to extract visual- and contact-based information from a single camera input. By placing a See-Thru-Network in front of a single in-finger camera, two layers of information can be obtained: one at a closer range for the soft interface, awaiting contact interactions, and another at a distance but occluded, about the scene. Contact-based perception is substituted markerlessly by reasoning about the masked deformation, facilitating reactive see-through grasping.
@@
Several directions remain open. Promising avenues include underwater perception and manipulation (Khatib et al., 2016), embodied robot learning with multimodal feedback that integrates SPN’s omni-directional adaptability and precise 6D force/torque sensing (Chi et al., 2024), and extending the single-camera approach to stereo vision for improved 3D reconstruction or 6D pose estimation.
@@ Acknowledgement
This work was partly supported by the National Natural Science Foundation of China under Grant 62473189.
Data Availability
All data and codes related to this work are hosted on the following project website: https://github.com/ancorasir/SeeThruFinger.